# Self-Routing RAG: Binding Selective Retrieval with Knowledge Verbalization

### v2

@@ Abstract
Selective retrieval improves the accuracy and efficiency of retrieval-augmented generation (RAG) by reducing distractions from low-quality retrievals. However, existing approaches underutilize the inherent knowledge of large language models (LLMs), leading to suboptimal retrieval decisions and degraded generation performance. To bridge this gap, we propose Self-Routing RAG (SR-RAG), a novel framework that binds selective retrieval with knowledge verbalization. SR-RAG enables an LLM to dynamically decide whether to retrieve external knowledge or verbalize its own parametric knowledge. To this end, we design a multi-task objective that jointly optimizes an LLM for knowledge source selection, knowledge verbalization, and response generation. SR-RAG further incorporates a nearest neighbor search mechanism at inference time to improve the accuracy of knowledge source decisions under domain shifts. Fine-tuning three LLMs with SR-RAG significantly improves both their response accuracy and reduces the inference latency. Compared to the strongest selective retrieval baseline, SR-RAG reduces the number of retrievals by 29% while improving performance by 5.1%.
@@ 1 Introduction
Retrieval-augmented generation (RAG) equips large language models (LLMs) with external knowledge sources at inference time, enabling stronger performance on tasks requiring up-to-date or domain-specific information (Khandelwal et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022; Ram et al., 2023; Shi et al., 2024). Recently, selective retrieval—an inference strategy that avoids unnecessary retrieval augmentations—has shown promising results in reducing distractions from low-quality retrievals and improving the efficiency of RAG (He et al., 2021; Mallen et al., 2023; Xu et al., 2024; Wu et al., 2024a).
However, a core question has been overlooked by current selective retrieval research:
Does selective retrieval fully honor the knowledge embedded in the LLM itself?
When retrieval is skipped, existing methods use a standard yet simplistic fallback: letting the LLM directly generate the response (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024; Wu et al., 2024a). This design prevents the LLM from explicitly articulating its parametric knowledge before generating a response. We argue that the ability to verbalize knowledge, while subtle, has a critical impact on the success of selective retrieval. First, verbalizing internal knowledge expands the LLM’s capacity to answer without retrieval. Prior work demonstrates the ability of LLMs to directly generate high-quality knowledge (Yu et al., 2023) as well as intermediate reasoning paths (Wei et al., 2022; Allen-Zhu and Li, 2025) to benefit the system’s performance. This is especially valuable for complex queries, where naive retrieval methods return irrelevant results and compute-intensive retrieval can return noisy contexts. Second, knowledge verbalization enables more accurate selective retrieval decisions. Existing works train retrieval policies by comparing RAG with direct LLM responses (Wang et al., 2023; Wu et al., 2024a) or resorting to likelihood-based preferences (He et al., 2021; Xu et al., 2024). By contrast, through explicit knowledge elicitation, knowledge verbalization helps characterize the LLM’s capabilities more precisely. Therefore, we argue that knowledge verbalization is a core component of selective retrieval to embrace.
Figure 1: An overview of SR-RAG. Given a user query, the system first selects the most appropriate knowledge source by combining special token prediction with nearest neighbor search. Then, the knowledge is either retrieved from an external source or self-verbalized by the LLM. Finally, the LLM forms the response based on the query and the knowledge. All the steps are streamlined into a single left-to-right generation pass.
We propose Self-Routing RAG (SR-RAG), a selective retrieval framework that tightly integrates knowledge verbalization. By reformulating selective retrieval as a knowledge source selection problem, SR-RAG enables an LLM to self-route between retrieving external knowledge and verbalizing its own parametric knowledge, as illustrated in Figure˜1. Recognizing the limitations of existing frameworks that fine-tune LLMs for selective retrieval via special token prediction (Asai et al., 2024; Wu et al., 2024a), SR-RAG introduces three key innovations. First, we generate diverse verbalizations of the LLM’s internal knowledge to create more accurate training labels for knowledge source selection (Section˜3.3). Second, we introduce a multi-task alignment objective that couples source selection, verbalization, and response generation. We further leverage self-supervised preference alignment over variants of verbalized knowledge to promote high-quality knowledge generation (Section˜3.3). Finally, existing approaches suffer from poor source decision accuracy at inference time due to domain shifts and LLM ability shifts caused by fine-tuning. To bridge this gap, SR-RAG proposes dynamic knowledge source inference via nearest neighbor search, augmenting likelihood-based retrieval decisions with neighboring policy examples in the hidden representation space of the fine-tuned LLM (Section˜3.4). Crucially, SR-RAG’s inference remains efficient, requiring only a single left-to-right generation pass.
@@
We evaluate SR-RAG by fine-tuning Llama-2-7B-Chat (Touvron et al., 2023), Phi-3.5-Mini-Instruct (Abdin et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024) on a mixture of knowledge-intensive tasks to develop the LLMs’ ability to select knowledge sources and improve the quality of knowledge verbalization. Extensive experiments across four benchmarks demonstrate that SR-RAG greatly outperforms both always retrieving and the baseline selective retrieval approach. Compared to the latter, SR-RAG achieves 8.5%/2.1%/4.7% higher overall performance while performing 26%/40%/21% fewer retrievals across the three LLMs respectively (Section˜5.2). Further analyses reveal that SR-RAG improves both the accuracy of selective retrieval decisions (Section˜5.3) and the overall inference efficiency (Section˜5.4). Finally, we conduct thorough ablation studies that confirm the necessity of all three core components to SR-RAG’s strong performance (Section˜5.5). Data and code will be publicly released at https://github.com/xiaowu0162/self-routing-rag.
@@ 2 Related Work
To enhance the efficiency of RAG systems and avoid potentially harmful retrievals, several works have proposed selectively skipping retrieval augmentation, which we call selective retrieval following Xu et al. (2024) and Wu et al. (2024a). One popular approach is to assess whether retrieval augmentation increases the likelihood of the LLM generating the correct answer and distill this observation into a supervised decision model (He et al., 2021; Schick et al., 2023; Xu et al., 2024). Analogously, Wang et al. (2023) and Wu et al. (2024a) directly evaluate the correctness of the answers generated with and without retrieval to create the supervision signal. Recent works have also explored solely examining the question’s nature to judge the need for retrieval on question answering tasks (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024). To further incorporate the LLM’s confidence in the retrieval decision, Ding et al. (2024), Yao et al. (2024), and Moskvoretskii et al. (2025a) use uncertainty-based selective retrieval. By contrast, this paper highlights the benefits of incorporating knowledge verbalization to make precise selective retrieval decisions and to boost the performance when retrieval is skipped.
@@
This work also relates to the broader field of adaptive RAG, which aims to develop configurable and instance-specific RAG inference strategies. Prior work has introduced active retrieval, where the system dynamically refines or re-issues queries if the initially retrieved content is insufficient (Jiang et al., 2023; Su et al., 2024). Other lines of research explore query decomposition and iterative retrieval to better handle complex questions by breaking them down into manageable sub-queries (Shao et al., 2023; Kim et al., 2023; Liu et al., 2024; Lee et al., 2024). Given the retrieval results, Asai et al. (2024) and Yan et al. (2024) propose inference strategies that critique or revise the retrieved knowledge to improve output quality. Parekh et al. (2025) incorporate an initial decision step to adaptively select the most suitable strategy based on the question. While our work focuses on the more targeted problem of selective retrieval, we introduce a novel form of adaptivity: enabling the LLM to self-route between external retrieval and internal knowledge verbalization, enhancing both flexibility and efficiency in a principled way.
@@
A growing body of work has explored using LLMs to generate auxiliary knowledge. Early studies demonstrated the potential of LLMs to produce relevant background knowledge in zero-shot settings for commonsense reasoning (Shwartz et al., 2020; Liu et al., 2022). More broadly, Yu et al. (2023) introduced a generate-then-read approach, treating LLMs as context generators to replace external retrieval for RAG. In addition, LLMs have been shown to generate effective intermediate reasoning steps for complex reasoning questions (Wei et al., 2022; Kojima et al., 2022). Building on these insights, this paper explores fully leveraging LLMs’ knowledge generation ability to benefit selective retrieval systems.
@@ 3 Approach
In this section, we first reformulate selective retrieval as a knowledge source selection problem. Then, we introduce the details of the proposed Self-Routing RAG framework.
@@ 3.1 Problem Formulation
Given a user query $q$, a knowledge source $s$ is invoked to return relevant information as a text sequence $s(q)$, which is then consumed by an LLM reader $M$ to generate a response $M(q,s(q))$. Standard RAG follows this paradigm, where $s$ typically represents a retriever over an external datastore.
@@
In practice, multiple knowledge sources may be available, denoted as $\mathcal{S}=\{\phi,s_{1},...,s_{N}\}$, where $\phi$ is a null knowledge source that returns nothing for any query. A knowledge source selector $P$ chooses the most appropriate source for a given query, i.e., $P(q,\mathcal{S})\in\mathcal{S}$. Selective retrieval can be viewed as the special case where $\mathcal{S}=\{\phi,s\}$ with the full pipeline expressed as $M(q,P(q,\mathcal{S})(q))$.
@@ 3.2 Self-Routing RAG: Overview
We propose Self-Routing RAG (SR-RAG), a generalization of selective retrieval that treats the LLM itself as a first-class knowledge source. As illustrated in Figure˜1, given a query $q$, the LLM self-determines which knowledge source to use: either retrieving from an external source or verbalizing its parametric knowledge. The final response is then generated based on both the query and the knowledge collected from the selected source.
Figure 2: Compared to traditional selective RAG, SR-RAG enables an LLM to self-route between knowledge sources and self-act as a knowledge source. We use blue to represent external information and red to represent the LLM and its self-generated tokens.
Building upon traditional selective retrieval methods (Asai et al., 2024; Wu et al., 2024a), SR-RAG fine-tunes the LLM to streamline its inference process with special tokens, enabling efficient inference with a single left-to-right generation pass. Three sets of special tokens are introduced:
@@
    <EOQ> marks the end of the query and prompts the LLM for knowledge source selection [^1] This design aligns with Wu et al. (2024a) but diverges from Asai et al. (2024). We argue that this token is necessary for the LLM to allocate probability mass to the tokens for the knowledge sources..
@@
    A set of special tokens <s>, each representing a knowledge source $s$. Our main setup uses two sources: external ($S_{e}$) and internal ($S_{i}$). This formulation can naturally accommodate more than two knowledge sources as well.
@@
    <EOK> indicates the end of the knowledge and triggers the LLM to generate the answer.
As shown in Figure˜2, compared to traditional selective retrieval, SR-RAG enables the LLM to actively select the best knowledge source and seamlessly act as a knowledge source itself.
@@ 3.3 Self-Routing RAG: Training
To train the backbone LLM for SR-RAG, we propose a pipeline to mine self-supervision from widely available question answering or instruction-following data, consisting of pairs of question $q$ and response $a$. Our pipeline uses the LLM itself as the internal knowledge source $S_{i}$, alongside the external knowledge source $S_{e}$, without requiring additional human supervision or synthetic labels from stronger LLMs, demonstrating strong scalability.
@@
To enable an LLM to accurately determine whether a question falls within its parametric knowledge and to robustly elicit the knowledge, we argue that thorough and diverse knowledge verbalization is crucial. Following this intuition, SR-RAG collects contexts by exploring each of the available knowledge sources:
  • Parametric Knowledge Verbalization: We leverage GenRead (Yu et al., 2023) to elicit knowledge from the LLM parametric knowledge source $S_{i}$ and generate $n$ diverse verbalized contexts, denoted as $c_{i_{1}},c_{i_{2}},...,c_{i_{n}}$.
  • External Knowledge Retrieval: We retrieve $n$ context chunks from the external knowledge source $S_{e}$, denoted as $c_{e_{1}},c_{e_{2}},...,c_{e_{n}}$. In this work, we consider retrieving from Wikipedia with an off-the-shelf dense retriever.
Then, each context $c_{j}$ is scored using the log-likelihood $l_{j}=p_{M}(a|q,c_{j})$ of generating the correct answer. Based on the ranking of $l_{j}$, we identify the preferred source $s\in\{S_{i},S_{e}\}$ as the one contributing to the majority of the top-$n$ ranked contexts. The resulting $(q,a,s,\{c_{j},l_{j}\})$ tuples are saved for model training. For convenience of later reference, we denote the contexts from $S_{i}$ and $S_{e}$ that lead to the highest and lowest likelihoods as $c_{i+},c_{i-},c_{e+},c_{e-}$, respectively. We present the formal data creation algorithm in Section˜B.1.
@@
SR-RAG proposes a two-stage multi-task learning framework that jointly optimizes knowledge source selection, knowledge verbalization, and response generation. The first stage performs behavior cloning on three losses:
@@
    $\mathcal{L}_{src}$: a cross-entropy loss for the preferred knowledge source $s$ following <EOQ>:
    $\mathcal{L}_{src}=-\log p_{M}(\texttt{\textless s\textgreater}|q),$ (1)
    where <s> represents the actual token corresponding to the chosen source $s\in\mathcal{S}$.
@@
    $\mathcal{L}_{verb}$: a cross-entropy loss on the knowledge tokens, only when the LLM itself ($S_{i}$) is labeled as the preferred knowledge source:
    $\mathcal{L}_{verb}=\begin{cases}-\log p_{M}(c_{i+}|q),&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$ (2)
@@
    $\mathcal{L}_{ans}$: a cross-entropy loss on generating the answer based on the query and the knowledge from the preferred source:
    $\mathcal{L}_{ans}=\begin{cases}-\log p_{M}(a|q,c_{i+}),&\text{if }s=S_{i},\\ -\log p_{M}(a|q,c_{e+}),&\text{if }s=S_{e}.\end{cases}$ (3)
The final loss for the first stage is a simple combination of the three objectives:
@@
To further boost the LLM’s ability to generate useful knowledge, SR-RAG incorporates a second-stage fine-tuning via direct preference optimization (DPO) (Rafailov et al., 2023), pairing self-verbalized knowledge with self-generated preference labels $(c_{i+},c_{i-})$.
@@
$\mathcal{L}_{verb}^{DPO}=\begin{cases}-\log\sigma\Big(\beta\log\frac{p_{M}(c_{i+}|q)}{p_{ref}(c_{i+}|q)}-\beta\log\frac{p_{M}(c_{i-}|q)}{p_{ref}(c_{i-}|q)}\Big),&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$
$M$ and refref are initialized with the LLM fine-tuned on $\mathcal{L}_{stage1}$, and only $M$ is updated.
Overall, this self-supervised pipeline effectively binds knowledge verbalization with the selective retrieval paradigm, enabling the LLM to learn accurate knowledge source preferences through performance-oriented labeling. Analogous to distilling complex “System 2” reasoning into fast “System 1” inference (Yu et al., 2024), the DPO objective leverages the computationally expensive high-quality knowledge to teach the LLM cost-efficient knowledge verbalization at inference time. Finally, SR-RAG naturally extends to more than two knowledge sources (further discussions and experiments in Appendix˜C), which is useful for distinguishing domain-specific corpora or retrieval methods with varying cost-quality trade-offs.

### v3

@@ Abstract
Selective retrieval aims to make retrieval-augmented generation (RAG) more efficient and reliable by skipping retrieval when an LLM’s parametric knowledge suffices. Despite promising results, existing methods are constrained by a binary design choice: either retrieve from a single external source or skip retrieval and let the LLM directly produce the final answer. We argue that this fallback underestimates the model’s knowledge and obscures the more general multi-source decision problem that arises in practical systems. We propose Self-Routing RAG (SR-RAG), which casts selective retrieval as knowledge source selection and treats the LLM itself as a first-class knowledge source. SR-RAG learns to select an appropriate knowledge source, optionally verbalize parametric knowledge, and answer using the selected source, all within a single left-to-right generation pass. SR-RAG further augments source selection by combining LLM-based uncertainty with a flexible external policy datastore to improve decision calibration. Across four benchmarks and three 7B-class LLMs, SR-RAG outperforms a strong selective retrieval baseline by 8.5%/2.1%/4.7% while performing 26%/40%/21% fewer retrievals, and it achieves favorable accuracy-latency trade-offs without dataset-specific threshold tuning.
@@ 1 Introduction
Retrieval-augmented generation (RAG) equips large language models (LLMs) with external knowledge at inference time, improving performance on tasks that require domain-specific or up-to-date knowledge (DBLP:conf/iclr/KhandelwalLJZL20; DBLP:conf/nips/LewisPPPKGKLYR020; DBLP:conf/icml/BorgeaudMHCRM0L22; ram-etal-2023-context; shi-etal-2024-replug). However, retrieval can be unnecessary when retrieved passages overlap with the LLM’s parametric knowledge, or even harmful when the passages are noisy or irrelevant. These limitations motivate selective retrieval, where the system retrieves only when retrieval is likely to be useful (he-etal-2021-efficient; mallen-etal-2023-trust; DBLP:conf/iclr/XuSC24; DBLP:conf/iclr/AsaiWWSH24; repoformer).
Figure 1: Self-Routing RAG Inference. (1) A knowledge source is selected based on a query by combining the LLM’s likelihood of a special source token with neighbor decisions from an external policy datastore. (2) If an external source is selected, the passages are retrieved and appended as context. If the internal source is selected, the LLM generates a compact verbalized knowledge context. (3) The LLM then produces the final answer conditioned on the query and the selected knowledge. All steps can be executed in a single left-to-right generation pass.
A limitation of current selective retrieval methods is reducing selective retrieval to a binary decision between consulting a single external source and a no-retrieval fallback where the LLM directly generates the final answer (mallen-etal-2023-trust; jeong-etal-2024-adaptive; DBLP:conf/iclr/AsaiWWSH24; repoformer). This design provides a coarse characterization of what the model can do without retrieval, as direct answer generation removes an opportunity to explicitly surface relevant parametric knowledge before committing to an answer. Consequently, routers trained by comparing RAG against direct answering, either via answer correctness (wang-etal-2023-self-knowledge; repoformer) or likelihood (he-etal-2021-efficient; DBLP:conf/iclr/XuSC24), can learn preferences inconsistent with the LLM’s non-retrieval capability. We observe this in a pilot study, where eliciting and conditioning on verbalized knowledge changes the preferred source for a substantial fraction of instances (Section˜5.1). Furthermore, the binary setup is inflexible as it cannot learn to leverage multiple knowledge sources, which could be an important limitation in real-world deployments.
We propose Self-Routing RAG (SR-RAG), which addresses both limitations by casting selective retrieval as a knowledge source selection problem and treating the LLM itself as a first-class knowledge source. SR-RAG expands the selective retrieval action space to routing between an internal knowledge source and an expandable set of external sources. As shown in Figure˜1, at inference time, when the internal source is selected, the LLM first verbalizes query-relevant parametric knowledge as a compact context, then answers conditioned on it. When an external source is selected, decoding is paused to retrieve evidence from the selected index, and the model answers conditioned on the retrieved passages.
To make SR-RAG reliable and efficient, we propose three techniques that span data construction, training, and inference. (1) Verbalization-aware supervision: given each training question, we collect multiple candidate contexts from each source, including diverse self-verbalized knowledge paragraphs, and assign the preferred source as the one whose contexts most increase the likelihood of the gold answer. This produces source selection labels that better reflect the LLM’s capability without retrieval than direct-answer baselines. (2) Multi-task learning: we jointly train an LLM as a better knowledge source router, a better knowledge source, and a better knowledge reader in RAG. A two-staged multi-task objective is proposed that couples source selection, knowledge verbalization, and answer generation, including a preference alignment training to further boost the knowledge verbalization ability. (3) kNN-augmented source selection: purely relying on LLM-based uncertainty to select a source as done in previous work is brittle and agnostic to training-induced LLM ability shifts. We propose to build a policy datastore from light rollouts and combine the model’s source selection probabilities with neighbor-defined source distributions in the hidden state space at test time.
@@
We evaluate SR-RAG by fine-tuning Llama-2-7B-Chat (touvron2023llama), Phi-3.5-Mini-Instruct (abdin2024phi), and Qwen2.5-7B-Instruct (yang2024qwen2) on a mixture of knowledge-intensive tasks and testing on four benchmarks. SR-RAG outperforms a strong standard selective retrieval baseline by 8.5%/2.1%/4.7% while performing 26%/40%/21% fewer retrievals across the three LLMs, respectively (Section˜5.2). We further show that SR-RAG makes more accurate source selection decisions (Section˜5.3) and achieves favorable accuracy-latency trade-offs without dataset-specific threshold tuning (Section˜5.4, Figure˜4). Ablations confirm that verbalization-aware labeling, preference alignment, and kNN-augmented inference each contribute materially to these gains (Section˜5.5). Finally, we experiment on routing among three sources (LLM, Wikipedia, and PubMed), where SR-RAG learns to retrieve from specialized corpus when needed while retaining strong performance on general workloads (Appendix˜B). Data and code will be publicly released at https://github.com/xiaowu0162/self-routing-rag.
@@ 2 Related Work
To enhance the efficiency of RAG systems and reduce harmful or distracting retrievals, several works propose selectively skipping retrieval augmentation (he-etal-2021-efficient; mallen-etal-2023-trust; DBLP:conf/iclr/XuSC24; repoformer). A common approach assesses whether retrieval increases the likelihood of generating the correct answer and distills this signal into a supervised decision model (he-etal-2021-efficient; schick2024toolformer; DBLP:conf/iclr/XuSC24). Similarly, wang-etal-2023-self-knowledge and repoformer compare answers generated with and without retrieval to form supervision based on correctness. Other lines of work examine the question alone to decide whether retrieval is needed (mallen-etal-2023-trust; jeong-etal-2024-adaptive; DBLP:conf/iclr/AsaiWWSH24). To incorporate model confidence, recent methods use uncertainty-based selective retrieval (DBLP:journals/corr/abs-2402-10612; DBLP:journals/corr/abs-2406-19215; moskvoretskii2025adaptiveretrievalselfknowledgebringing). In contrast to prior work that treats skipping retrieval as direct answer generation, our work highlights that routing decisions and non-retrieval performance both improve when the LLM is treated as a knowledge source and is trained to verbalize query-relevant parametric knowledge.
@@
This work also relates to adaptive RAG, which studies instance-specific inference strategies beyond a fixed retrieve-then-read pipeline. Prior work studies active retrieval, where the system re-issues queries when initial retrieval is insufficient (jiang-etal-2023-active; su-etal-2024-dragin). Another line explores query decomposition and iterative retrieval to handle complex questions by breaking them into sub-queries (shao-etal-2023-enhancing; kim-etal-2023-tree; liu-etal-2024-ra; lee-etal-2024-planrag). Given retrieved results, DBLP:conf/iclr/AsaiWWSH24 and DBLP:journals/corr/abs-2401-15884 critique or revise the retrieved knowledge to improve output quality. parekh2024dynamic incorporate an initial decision step to adaptively select the most suitable strategy based on the question. While SR-RAG focuses on the targeted decision of which knowledge source to consult, it introduces a complementary form of adaptivity by allowing the model to route among heterogeneous sources, including its own parametric knowledge, within a single generation pass.
@@
A growing body of work explores using LLMs to generate auxiliary knowledge. Early studies show that LLMs can produce relevant background knowledge in zero-shot settings for commonsense reasoning (shwartz-etal-2020-unsupervised; liu-etal-2022-generated). More broadly, DBLP:conf/iclr/0002IWXJ000023 propose a generate-then-read approach, using the LLM as a context generator in place of external retrieval for RAG. LLMs have also been shown to generate effective intermediate reasoning steps for complex questions (DBLP:conf/nips/Wei0SBIXCLZ22; DBLP:conf/nips/KojimaGRMI22). Building on these insights, our work integrates LLM-generated knowledge into selective retrieval by making it an explicit, routable source, and by training the model to both decide when to rely on it and to produce compact verbalizations that improve end-to-end answering.
@@ 3 Approach
We first reformulate selective retrieval as knowledge source selection. We then present Self-Routing RAG (SR-RAG), a framework that allows an LLM to self-route between external retrieval and internal knowledge verbalization within a single left-to-right generation pass.
@@ 3.1 Problem Formulation
Given a user query $q$, a knowledge source $s$ returns relevant information as a text sequence $s(q)$ (possibly empty). An LLM reader $M$ then produces a response conditioned on both the query and the returned knowledge:
$M(q,s(q)).$ (1)
Under this framework, standard RAG is an instantiation with $s$ being a retriever over an external datastore and $s(q)$ being retrieved passages.
@@
In many practical settings, multiple knowledge sources may be available, denoted by
$\mathcal{S}=\{\phi,s_{1},\ldots,s_{N}\},$ (2)
where $\phi$ is a null source that always returns an empty string. A knowledge source selector $P$ chooses the most appropriate source for a given query, i.e., $P(q,\mathcal{S})\in\mathcal{S}$. Combined with the downstream inference, the full pipeline becomes
$M\big(q,P(q,\mathcal{S})(q)\big).$ (3)
Selective retrieval corresponds to the binary case $\mathcal{S}=\{\phi,s\}$, where the selector decides whether to invoke the retriever $s$ or to skip retrieval via $\phi$.
@@ 3.2 Self-Routing RAG: Overview
We propose Self-Routing RAG (SR-RAG), which generalizes selective retrieval by treating the LLM itself as a first-class knowledge source. Concretely, SR-RAG routes between external sources $\{S_{e1},...,S_{em}\}$ (e.g., several external corpora) and an internal source $S_{i}$ (parametric knowledge). If $S_{i}$ is selected, the model does not directly answer. Instead, it first verbalizes query-relevant parametric knowledge into a compact context and then answers conditioned on this context. This makes the non-retrieval path an explicit knowledge consultation step rather than an absence of context.
Figure 2: SR-RAG expands selective retrieval into source routing. Compared to traditional selective retrieval, SR-RAG allows the LLM to route among knowledge sources and to act as a knowledge source itself. Blue denotes external retrieved content and red denotes the LLM and its self-generated tokens.
Following prior work that streamlines adaptive RAG inference via special-token prediction (DBLP:conf/iclr/AsaiWWSH24; repoformer), SR-RAG uses three sets of special tokens to implement routing in a single left-to-right generation pass:
@@
    <EOQ>, which marks the end of the query and prompts source selection. [^1] This design aligns with repoformer but differs from DBLP:conf/iclr/AsaiWWSH24. We find the additional <EOQ> helps the model allocate probability mass to source tokens without interfering with the query content.
@@
    A set of source tokens <s>, each representing a knowledge source $s\in\mathcal{S}$. [^2] Our main setup uses a simple Wikipedia-based retrieval <Wiki> as the only $S_{e}$ and <Self> for $S_{i}$. This formulation naturally extends to more than two sources by adding additional source tokens. We study a three-source setting in Appendix B.
@@
    <EOK>, which indicates the end of the knowledge context and triggers answer generation.
An illustration of the difference between SR-RAG and standard selective RAG is provided in Figure˜2.
@@ 3.3 Self-Routing RAG: Training
We propose a fine-tuning recipe to improve any LLM’s ability as the core model in the SR-RAG framework. Inspired by repoformer, we propose to mine self-supervision from existing question-answer data $(q,a)$. Our pipeline uses the LLM itself as the internal source $S_{i}$ along with the external knowledge source, requiring neither human routing labels nor a stronger teacher model. For simplicity, we formulate with $\mathcal{S}=\{S_{i},S_{e}\},$ here. The more general form readily follows and is formulated in Appendix˜B.
@@
Our central goals are (1) labeling which source is more helpful for a given instance (2) collecting the most helpful knowledge when parametric knowledge is preferred. To do so, our rollout process first collects $n$ candidate contexts from each of the knowledge sources:
  • Parametric knowledge verbalization: We use GenRead (DBLP:conf/iclr/0002IWXJ000023) to elicit $n$ diverse verbalizations from the internal source $S_{i}$, denoted as $c_{i_{1}},\ldots,c_{i_{n}}$.
  • External knowledge retrieval: We retrieve $n$ context passages from the external source $S_{e}$, denoted as $c_{e_{1}},\ldots,c_{e_{n}}$.
Then, each candidate context $c_{j}$ is scored by the log-likelihood of generating the gold answer:
$l_{j}=\log p_{M}(a\mid q,c_{j}).$ (4)
We then rank all $2n$ contexts by $l_{j}$ and select the preferred source $s\in\{S_{i},S_{e}\}$ as the source that appears most frequently among the top-ranked contexts. For convenience, we denote the highest- and lowest-scoring contexts from each source as $c_{i+},c_{i-},c_{e+},c_{e-}$, respectively. The resulting training tuple is $(q,a,s,\{c_{j},l_{j}\})$. The formal data collection algorithm is provided in Section˜A.2.
@@
SR-RAG fine-tunes the LLM to perform (i) source selection, (ii) knowledge verbalization when $S_{i}$ is selected, and (iii) answer generation conditioned on the selected knowledge. We use a two-stage training procedure.
Stage 1 We perform behavior cloning by jointly optimizing three losses:
@@
    $\mathcal{L}_{src}$, a cross-entropy loss that trains the model to predict the preferred source token immediately after <EOQ>:
    $\mathcal{L}_{src}=-\log p_{M}(\texttt{\textless s\textgreater}\mid q).$ (5)
@@
    $\mathcal{L}_{verb}$, a cross-entropy loss over the knowledge tokens when the LLM itself is labeled as the preferred knowledge source:
    $\mathcal{L}_{verb}=\begin{cases}-\log p_{M}(c_{i+}\mid q),&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$ (6)
@@
    $\mathcal{L}_{ans}$, a cross-entropy loss on generating the answer conditioned on the preferred knowledge:
    $\mathcal{L}_{ans}=\begin{cases}-\log p_{M}(a\mid q,c_{i+}),&\text{if }s=S_{i},\\ -\log p_{M}(a\mid q,c_{e+}),&\text{if }s=S_{e}.\end{cases}$ (7)
We combine these objectives as
@@
Stage 2 To further improve the knowledge verbalization quality, we apply Direct Preference Optimization (DPO, DBLP:conf/nips/RafailovSMMEF23) using self-generated preference pairs $(c_{i+},c_{i-})$:
@@
$\mathcal{L}_{verb}^{DPO}=\begin{cases}-\log\sigma\Big(\beta\log\frac{p_{M}(c_{i+}\mid q)}{p_{ref}(c_{i+}\mid q)}-\beta\log\frac{p_{M}(c_{i-}\mid q)}{p_{ref}(c_{i-}\mid q)}\Big),&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$
Here $M$ and refref are initialized from the stage 1 model checkpoint, and only $M$ is updated. Conceptually, stage 2 distills a compute-intensive “system 2” process that searches over multiple verbalizations into a fast “system 1” behavior that produces a single high-quality knowledge verbalization during standard left-to-right decoding.